mix a solution that is 50% acid with a solution that is 100% water to make 4L of a solution that is 10% acid. How much of each solution should you use?
Answer by TimothyLamb(4379)

x = liters of 50% acid = liters of 50% water
y = liters of pure water
z = liters of 10% acid = liters of 90% water = 4
---
x + y = 4
50x + 100y = 90(x + y)
50x + 100y = 90*4
---
put the system of linear equations into standard form
---
x + y = 4
50x + 100y = 360
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x = liters of 50% acid = 0.8
y = liters of pure water = 3.2
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php